OpenStudy (anonymous):

Find a second-degree polynomial P such that P(2) = 3, P'(2) = 7, and P''(2) = 6.

OpenStudy (anonymous):

\[f \left( x \right)=ax^2+bx+c \text{, }f'\left( x \right)=2ax+b \text{, }f''\left( x \right)=2a\] plug in and solve

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

what did you get?